OpenStudy (anonymous):

Choose the correct x-intercept and y-intercept of the equation 3x − 2y = 6. x-intercept: (0, −3), y-intercept: (0, 2) x-intercept: (−3, 0), y-intercept: (2, 0) x-intercept: (0, 2), y-intercept: (0, −3) x-intercept: (2, 0), y-intercept: (0, −3)

OpenStudy (anonymous):

Find the y- intercept and the x intercept for the equation. -15x + y = 55 so I am looking for y=mx+b? if so y= 15x + 55 y x 55 = 15x/55 the answer is y=15/55x + 55?